f(x)=-16x^2+3600 solve for x
Found 3 solutions by mananth, Alan3354, ikleyn:
Answer by mananth(16946)

f(x)=-16x^2+3600 solve for x
f(x)=-16x^2+3600
equate to 0
3600-16x^2=0
(60)^2-(4x)^2=
Difference of two squares
a^2-b^2= (a+b)(a-b)
(60+4x)(60-4x)=0
60=-4x or 60 =4x
x=-15 Or x=15



Answer by Alan3354(69443)

f(x)=-16x^2+3600 solve for x
--------------
No value given for f(x).
x is dependent on f(x), or v.v., no reason to assume it's zero.
------------
If it's "Find the zeros," then it can be solved.

Answer by ikleyn(52750)

.
f(x)= -16x^2+3600 solve for x
~~~~~~~~~~~~~~~~~~~


Mathematically, it is INCORRECT and NON-sensical assignment.

It is non-sensical, because it makes no sense.


A person who created it as a Math assignment, is mathematically ILLITERATE.

It is mathematically non-sensical and therefore can not be considered;
should not be considered, and should not even be discussed.

The right place for such compositions is the closest garbage bin.


The closest meaningful request is "find the inverse function for f(x) = -16x^2 + 3600".

But in Math, it is not a right approach to hypothesize what the composer wanted to say in his composition.

Therefore, I will not develop further this my idea.


Mathematicians simply sort the upcoming statements in two categories.

They accept the mathematically correct statements/assignments for execution,
and they throw mathematically incorrect statements/assignments in garbage bin.


                    Not because they are so bad people.
            Simply because those are the rules of the game.